OpenStudy (anonymous):

Solving each system of linear equations using determinants... 1.) 3x-10y=15 2x+3y=-26 2.) 3x+2y=3 2x-5y=-17 3.) 4x-9y=2 5x+3y=-26 4.) 4x+3y=-9 2x+y=-5 5.) 2x+5y-22=0 x-3y=0 Please Please Solve. I need a good answer. Thankyou :) GBY.

OpenStudy (anonymous):

Answeeeeerr :(

OpenStudy (anonymous):

Maybe people don't want to answer u because you are asking 5 questions involving matrices which will take a lot of time for someone to answer. Besides, these are not difficult, just use Ax = b and x = A^-1b where A is the inverse matrix and do the calculations. Or maybe just post 1 of the questions and someone will maybe explain u how to do it, then you can do the others by yourself.

OpenStudy (anonymous):

ick this are annoying.

OpenStudy (anonymous):

first one is \[(-\frac{215}{29},-\frac{108}{29})\]

OpenStudy (anonymous):

you need the method?

OpenStudy (anonymous):

you here?

OpenStudy (anonymous):

says use determinants. the denominate of these is the determinant of 3 -10 2 3 namely \[3\times 3-(-10)\times 2=29\]

OpenStudy (anonymous):

denominator is the word i was looking for

OpenStudy (anonymous):

He replied in the other one saying he needs the method.

OpenStudy (anonymous):

jeez ok maybe i write it here. just was trying to say that to find y you take the determinant of 3 15 2 -26 get \[3\times -26-2\times 15=-108\]

OpenStudy (anonymous):

then put \[\frac{-108}{29}\] and similarly for x find the determinant of -10 26 15 3 namely \[-10\times 3-15\times 26=-215\]

OpenStudy (anonymous):

same for others. routine donkey work. i would use software